### v2

# Blade: A package for block-triangular form improved Feynman integrals decomposition

Abstract
In this article, we present the package Blade as the first implementation of the block-triangular form improved Feynman integral reduction method. The block-triangular form has orders of magnitude fewer equations compared to the plain integration-by-parts system, allowing for strictly block-by-block solutions. This results in faster evaluations and reduced resource consumption. We elucidate the algorithms involved in obtaining the block-triangular form along with their implementations. Additionally, we introduce novel algorithms for finding the canonical form and symmetry relations of Feynman integrals, as well as for performing spanning-sector reduction. Our benchmarks for various state-of-the-art problems demonstrate that Blade is remarkably competitive among existing reduction tools. Furthermore, the Blade package offers several distinctive features, including support for complex kinematic variables or masses, user-defined Feynman prescriptions for each propagator, and general integrands.
PROGRAM SUMMARY
Program title: Blade
Developer’s repository link: https://gitee.com/multiloop-pku/blade
Licensing provisions: MIT
Programming language: Wolfram Mathematica 11.3 or higher
External routines/libraries used: Wolfram Mathematica [1], FiniteFlow [2]
Nature of problem: Automatically reducing dimensionally regularized Feynman integrals into linear combination of master integrals.
Solution method: The program implements recently proposed block-triangular form to significantly improve the reduction efficiency.
Restrictions: the CPU time and the available RAM
Acknowledgements
We thank Li-Hong Huang, Xiang Li, Bernhard Mistlberger, Tiziano Peraro, Ao Tan, Zi-Hao Wu, and Yang Zhang for many useful communications and discussions. We thank Huan-Yu Bi, Long Chen, Xiang Chen, Yuan-Hong Guo, Chuan-Qi He, Rui-Jun Huang, Xu-Hang Jiang, Xiang Li, Huai-Min Yu and Yu-Dong Zhang for providing numerous invaluable feedback on the package. The work was supported in part by the National Natural Science Foundation of China (Grants No. 12325503, No. 11975029), the National Key Research and Development Program of China under Contracts No. 2020YFA0406400, and the High-performance Computing Platform of Peking University. The research of XG was also supported by the United States Department of Energy, Contract DE-AC02-76SF00515. The research of XL was also supported by the ERC Starting Grant 804394 HipQCD and by the UK Science and Technology Facilities Council (STFC) under grant ST/T000864/1. JaxoDraw [89] was used to generate Feynman diagrams.
Appendix A Propagators
Feynman integrals in Blade are considered in Minkowski space, $l^{2}$, of the denominators instead of ‘Euclidean’ space, $-l^{2}$.
We associate to each denominator a number ‘pre’ to indicate Feynman prescription,
$\begin{cases}D_{j}(l_{i})+\mathrm{i}0,&\textrm{if}~{}pre=1,\\ D_{j}(l_{i}),&\textrm{if}~{}pre=0,\\ D_{j}(l_{i})-\mathrm{i}0,&\textrm{if}~{}pre=-1.\end{cases}$
‘pre’ is set to 1 by default. Symmetry is detected among denominators that have identical prescriptions. More details can be found in the “example/1_prescription” folder.
Propagators that can be written as $q^{2}-c$ or $q\cdot p-c$ is called standard propagators in Blade. Here $q$ is a linear combination of loop momenta and external momenta, and $p$ is a linear combination of external momenta. The current version of Blade will not detect symmetry relations if other general propagators are present.
Appendix B Discretized Symmetries
B.1 Point group of Feynman integral
Let us consider a Feynman integral with $P$ propagator denominators, where each denominator depends on a propagating momentum $q_{a}$, defined as
$q_{a}=\sum_{i=1}^{L}\alpha_{a}^{i}l_{i}+\sum_{j=1}^{E}\beta_{a}^{j}p_{j},\quad a=1,\ldots,P.$ (45)
[…]
Figure 7: The choices of spanning tree (blue thin lines) and loop base (red thick lines) in a 2-loop self-energy diagram
In Section 2.3, we highlighted that linear transformations of loop momenta enable us to relate any given Feynman integral to its canonical form, which we will define next. The only condition we have imposed so far on the canonical form is that each loop momentum $l_{i}$ needs to serve as the propagating momentum of one of the propagators, as much as possible. We refer to the propagator with propagating momentum $l_{i}$ as the $i$-th loop base. However, there can be different choices of loop bases, and even for a given choice of loop bases, there can be different ways to label each propagating momentum. Starting from a candidate for the canonical form, which means a choice of loop bases has already been made, we can generate many more candidates by applying the following operations:
  • 1.
    Reverse propagating momentum of each propagator:
    $q_{a}\to q_{a}^{\prime}=-q_{a},$ (46)
    which is just a matter of defining the orientation of the momentum. This results in $2^{P}$ choices.
  • 2.
    Reverse each loop momentum:
    $l_{i}\to l_{i}^{\prime}=-l_{i},$ (47)
    resulting in $2^{L}$ choices.
  • 3.
    Choose loop bases from different propagators. The number of choices equals the number of terms in the 1st Symanzik polynomial, denoted as #$U$. For example, the diagram in Fig. 7 with internal lines labeled as shown has the 1st Symanzik polynomial
    $U(x_{1},x_{2},x_{3},x_{4},x_{5})$ (48)
    $=$ $(x_{1}+x_{3})(x_{2}+x_{4})+(x_{1}+x_{2}+x_{3}+x_{4})x_{5}$
    $=$ $x_{1}x_{2}+x_{1}x_{4}+x_{3}x_{2}+x_{3}x_{4}+x_{1}x_{5}+x_{2}x_{5}$
    $+x_{3}x_{5}+x_{4}x_{5},$
    which has 8 terms. In Fig.7, the choices of loop bases in diagrams (a), (b) and (c) correspond to the $1$st term $x_{1}x_{2}$, 3rd term $x_{3}x_{2}$, and 5th term $x_{1}x_{5}$, respectively. Note that this counting is correct if we only choose double propagators as loop bases.
  • 4.
    Permutation of loop momenta:
    $l_{i}\to l_{i}^{\prime}=l_{\sigma(i)},$ (49)
    where $\sigma$ is any permutation operation applying to the $L$ loop momenta. This results in $L!$ choices.
Thus, the total number of candidates for the canonical form is $2^{P}\times 2^{L}\times\#U\times L!$, which is usually very large, making it impractical to enumerate each choice and then sort them. Fortunately, since these operations form a point group with a well-organized structure, we can find a representative candidate with much less effort.
B.2 Canonical form of Feynman integral
We assume the Feynman integral under consideration is already a candidate for the canonical form. Otherwise, we can simply choose any loop bases by selecting a term in the 1st Symanzik polynomial and transform the integral to a candidate form. Our algorithm to find the canonical form, which is a representative candidate among all possibilities, can be summarized in the following five steps:
1)
Orientations of propagating momenta.
For each propagator $q_{a}$ defined in Eq. (45), we use a vector $v_{a}$ to denote it as
$v_{a}=(\alpha_{a}^{1},\cdots,\alpha_{a}^{L};\beta_{a}^{1},\cdots,\beta_{a}^{E}).$
We sort the list $\{v_{a},-v_{a}\}$ in alphabetical order for each propagator and replace the vector $v_{a}$ by the last one of the sorting result. Then, if two propagators belong to the same branch, the first $L$ terms of their vectors must be the same, which enables us to classify propagators into branches.
Figure 8: Partial loop bases in a length-$2$ branch
2)
Representative propagator in each branch.
For each branch, find the most special propagator by sorting all of them with different momenta orientations. The special one will be the candidate for the loop base in this branch. Loop momenta are irrelevant at this step.
Usually, there is only one special propagator. However, for a length-2 branch with the same mass in both propagators, as shown in Fig. 8, two pairs of representations are identical (b1, b4 and b2, b3). So, in the sub-diagram, neither is more special, and we need to retain both propagators for later consideration.
3)
Loop bases.
Use the 1st Symanzik polynomial of the corresponding vacuum diagram, obtained by removing external legs, to determine all possible choices of loop bases. Sort the obtained expressions of the integrand for all possible choices to find a set of special choices. Note that we will not specify the label and orientation of loop momenta in this step, so only the total number of loop momenta and the absolute values of external momenta are relevant in the sorting procedure. For example, for an integral with three loops and two external momenta, the propagating momenta will be mapped to the following form during the sorting procedure:
$\begin{split}q_{1}&=l_{1}-l_{2}+l_{3}+p_{1}\to(3;1,0),\\ q_{2}&=l_{2}-l_{1}-p_{2}\to(2;0,1).\end{split}$ (50)
4)
Orientations of loop momenta.
Beginning from the special choices in the previous step, iterate over all orientations of loop momenta, and sort the obtained expressions to find a class of special choices. As we will not specify the labels of loop momenta in this step, propagating momenta will be mapped to the following form during the sorting procedure:
$\begin{split}q_{1}&=l_{1}-l_{2}+l_{3}+p_{1}\to(1,1,-1;1,0),\\ q_{2}&=l_{2}-l_{1}-p_{2}\to(1,-1;0,-1),\end{split}$ (51)
where the numbers before the semicolon are in an orderless manner, and the numbers after the semicolon are ordered.
5)
Permutations of loop momenta.
Beginning from the special choices in the previous step, permute all loop momenta and sort the obtained expressions to find a class of special choices. All information for loop momenta and external momenta is specified in this step. For example, we have the following mapping relation:
$\ \begin{split}q&=l_{1}-l_{2}+l_{3}+p_{1}\to(1,-1,1;1,0),\\ q&=l_{2}-l_{1}-p_{2}\to(-1,1;0,-1).\end{split}$ (52)
Finally, we obtain a canonical form for the considered integral, as well as several linear transformations of loop momenta that can transform the original form to the canonical form. The complexity is roughly of the order $O(L!)$, insensitive to the number of external momenta. Although this complexity can be further reduced by splitting the fifth step to more steps, it is usually not necessary for practice use.
B.3 Simplified symmetry relations
With the canonical form, we can easily check whether two sectors are the same after applying a certain linear transformation of loop momenta. Furthermore, as there can be several transformations that keep the canonical form unchanged, known as “automorphic” maps, they can generate linear relations of Feynman integrals (FIs) within the same sector. These are symmetry relations that will be used in FI reduction.
For any symmetry relation, the propagator denominators are permuted by $\sigma$ while the ISP undergoes a general linear transformation $D_{a}\to D_{a}^{\prime}=\sum_{b}A_{ab}D_{b}$. More explicitly, we have
$\left\{\begin{matrix}\displaystyle D_{1}&\to&D^{{}^{\prime}}_{1}&=&D_{\sigma(1)},\\ &&\ldots\ldots&&\\ D_{k}&\to&D^{{}^{\prime}}_{k}&=&D_{\sigma(k)},\\ D_{k+1}&\to&D^{{}^{\prime}}_{k+1}&=&\displaystyle\sum_{j=1}^{N}A_{k+1,j}D_{j},\\ &&\ldots\ldots&&\\ D_{N}&\to&D^{{}^{\prime}}_{N}&=&\displaystyle\sum_{j=1}^{N}A_{N,j}D_{j}.\end{matrix}\right.$ (53)
For an integral
$\int\mathrm{d}\mu_{L}\frac{D_{k+1}^{-\nu_{k+1}}\ldots D_{N}^{-\nu_{n}}}{D_{1}^{\nu_{1}}\ldots D_{k}^{\nu_{k}}},$ (54)
we can generate a symmetry relation by replacing $D_{a}$ with $\sum_{b}A_{ab}D_{b}$ for all values of $a$. However, if the rank $R=-\sum_{a=k+1}^{N}\nu_{a}$ is not very small, the length of this relation can be extremely long, scaling as $O(N^{R})$ for $N\sim 10$ in cutting-edge problems.
In fact, the above integral can be re-expressed by any of the following kinds of symmetry relations:
$\int\mathrm{d}\mu_{L}\frac{(D_{k+1}^{(l)})^{-\nu_{k+1}}\ldots(D_{N}^{(l)})^{-\nu_{N}}}{D_{1}^{\nu_{1}}\ldots D_{k}^{\nu_{k}}}$
$=$ $\int\mathrm{d}\mu_{L}\frac{(D_{k+1}^{(r)})^{-\nu_{k+1}}\ldots(D_{N}^{(r)})^{-\nu_{N}}}{D_{\sigma(1)}^{\nu_{\sigma(1)}}\ldots D_{\sigma(k)}^{\nu_{\sigma(k)}}},$ (55)
where $(D_{a}^{(l)},D_{a}^{(r)})$ can be set to either $(D_{a},\sum_{b}A_{ab}D_{b})$ or $(\sum_{b}A^{-1}_{ab}D_{b},D_{a})$. By choosing $(D_{a}^{(l)},D_{a}^{(r)})$ for each value of $a$ to minimize the number of terms of the equation, we can find the simplest symmetry relation related to the integral in Eq. (54). This will roughly reduce the complexity from $O(N^{R})$ to its square root, $O(\sqrt{N^{R}})$.